OpenStudy (anonymous):

Still not getting this. help! I tried using the substitution method, it's not making sense. The problem says: What is the solution of the system of equations? 3x + 2y + z = 7 5x + 5y + 4z = 3 3x + 2y + 3z =1 I get z = -3 (if that's right) then I'm lost

OpenStudy (mathlegend):

1st you need to take the 1st and 2nd equation.... choose to cancel out a variable. I will choose z. (-4)3x + 2y + z = 7 5x + 5y + 4z = 3 -12x - 8y - 4z = -28 5x + 5y + 4z = 3 ----------------- -7x -3y = -25 (One Equation to come back and use)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

btw, z =-3, so that is correct.

OpenStudy (anonymous):

ok that makes sense so far

OpenStudy (anonymous):

I originally cancelled out y first

OpenStudy (mathlegend):

ok I'm going to do this on paper first okay?

OpenStudy (anonymous):

ok

OpenStudy (mathlegend):

Okay good...

OpenStudy (mathlegend):

@tms622 I will do it your way. You were fine in your logic.

OpenStudy (anonymous):

I got lost after I figured out z. Y and X confused me or something.

OpenStudy (mathlegend):

(5)3x+2y+z=7 (-2)5x+5y+4z=3 --------------- 15x+10y+5z=35 -10x-10y-8z=-6 ---------------- 5x-3z=29 (One Equation) (2)5x+5y+4z=3 (-5)3x+2y+3z=1 ---------------- 10x+10y+8z=6 -15x-10y-15z=-5 ---------------- -5x-7z=1 (Another Equation) 5x-3z=29 -5x-7z=1 ---------- -10z=30 z = -3

OpenStudy (mathlegend):

Now.... since we have 'z' we must go back to an equation that has only TWO variables. So either... 5x-3z=29 OR -5x-7z=1

OpenStudy (anonymous):

ok

OpenStudy (mathlegend):

Lets choose the all positive one. Then substitute in z. 5x-3(-3)=29 5x + 9 = 29 5x = 20 x = 4

OpenStudy (mathlegend):

Now we have TWO variables solved for... z=-3 x= 4 Now we need an equation with 3 variables in it.. then substitute those two in there.

OpenStudy (anonymous):

ohhh ok then solve for y

OpenStudy (mathlegend):

So we can choose out of.... the originals :) 3x + 2y + z = 7 5x + 5y + 4z = 3 3x + 2y + 3z =1

OpenStudy (mathlegend):

3(4)+2y-3=7 12 + 2y - 3 = 7 9+2y = 7 2y = -2 y = -1

OpenStudy (anonymous):

yay Thats what I just got too. Thanks for all your help. That makes alot more sense now.

OpenStudy (mathlegend):

:)